# A Survey on Sequential Recommendation

### v1

@@ 5.4.1 Common Features Fusion
In this section, we mainly introduce some SR models that fuse some common features (i.e., position, category, brand ,etc.) and item IDs. The structure of these SR models is features-agnostic to some extent. ICAI-SR [3] leverages a heterogeneous graph to capture complex item-to-attribute relationships by inner attribute aggregation and attribute-to-item aggregation, and then passes aggregation embeddings through an entity sequential model. MLP4Rec [117] designs a fusion mechanism to capture sequential, cross-channel and cross-feature correlations. SMLP4Rec [118] leverages a layer normalization for a sequence mixing module, a feature mixing module and a channel mixing module simultaneously. Meanwhile, to enhance efficiency and effectiveness, SMLP4Rec leverages a parallel mode. RetaGNN [119] leverages a relational attentive GNN in which learnable weight matrices focus on various relations among users, items and attributes to have inductive and transferable properties. Then, RetaGNN [119] leverages Transformer to capture short-term and long-term temporal patterns.
In Transformer-based SR models, several advancements have been made in integrating common features with item IDs. FDSA [120] first fuses different features by a vanilla attention mechanism. Then, FDSA leverages separate self-attention blocks to independently learn feature-level and item-level representations. Finally, FDSA fuses the two kinds of preferences for recommendation. SASRecF [25] is a sample model to fuse common feature embeddings and item ID embeddings through a concatenation operation. This fusion method is considered invasive [4], similar to an addition operation and a gating addition operation. The specific process is shown in part (a) of Figure 8. This fusion method might compromise the original information in item embeddings, potentially leading to negative effects. To address this issue, NOVA [4] leverages a non-invasive method to enhance the learning of attention matrices. In NOVA, key and query matrices are derived by integrating different feature embeddings and item embeddings while value matrices are directly obtained from item embeddings. However, NOVA [4] still leverages integrated embeddings to learn key and value matrices. The compounded embedding space might introduce random disturbances and share identical gradients, which might degrade recommendation performance. DIF-SR [25] leverages decoupled attention to adaptively learn different features by flexible gradients. MSSR [121] captures the associations among an item ID sequence and each feature sequence. ASIF [122] can maintain semantic consistency between item IDs and features. Meanwhile, ASIF [122] can make full use of different kinds of features without compromising item embeddings, thanks to contrastive learning.
@@ 5.4.2 Context Information Fusion
Context information can reflect the specific environment where behaviors happen. SR models can provide more accurate recommendations by considering the context information. Bi-CASER [123] uses a context graph to represent context information, and then leverages bi-Transformer to learn users’ sequential behaviors. CRNNs [124] incorporates context information (e.g., interaction timestamps) by combining item embeddings with context embeddings to recommend personalized items. STAR [125] leverages stacked RNNs to model the sequential and context-aware information simultaneously. STEM [126] considers item behaviors and leverages Transformer-based architecture for exploiting spatial and temporal information.
@@ 5.4.3 Special Features Fusion
In this section, we mainly introduce SR models designed to fuse special features. These SR models are typically applied to situations where certain special features (e.g., social networks and timestamps) are available. STEN [26] introduces temporal information and a social network to capture users’ fine-grained dynamic interests in an event-level direct paradigm. DUVRec [127] constructs a timespan-aware sequence graph and an attribute-augmented graph to learn item-view and factor-view user representations, respectively. Recommendation performance is then improved by fusing the two representations. Based on features and labels, STARec [128] retrieves a target user’s historical behaviors through a search-based retriever. Then, STARec [128] leverages graph structure to enhance representation learning. SGCross [129] transfers knowledge from an auxiliary domain to a target domain from a personal view, a temporal view and a collaborative view. The knowledge encompasses personal preferences, dynamic preferences and collaborative preferences.
@@ 6.1 Multi-Modal SR
A user can be represented by the user’s profile. The profile may consist of user ’s demographic information or user ’s interacted items. An item can be denoted by an item ID embedding or a pre-trained modality embedding. Modality content features consist of text, images, videos, audio, text-image multi-modal pairs, and so on. ID-agnostic representations can be learned using pre-trained modality models. A common text encoder is BERT, which is usually used to extract semantic features from text. Swin Transformer is an image encoder, which can be leveraged to extract features from images. The modal-based SR models are shown in Figure 9. Different from ID-based SR models, one type of modal-based SR models leverages a single modality feature (e.g., text or images) to leverage items. The other type of modal-based SR models leverages some modal features (e.g., text and images) to recommend personalized items [130]. In order words, the representations of item are obtained from the multi-modal features of item in modal-based SR models, which is formalized as follows:
@@
UniSRec [8] leverages an MoE-enhanced adaptor and parametric whitening for domain adaptation. To learn universal sequence representations better, UniSRec proposes a seq-item contrastive task and a seq-to-seq contrastive task. The model is pre-trained in one domain and subsequently fine-tuned in another domain using either an inductive or transductive approach. Meanwhile, UniSRec adopts two augmentation strategies (i.e., item drop and word drop). VQ-Rec [27] is a novel approach to learn vector-quantized item representations for solving the problem about over-emphasizing the effect of text features, which might lead to poor recommendation performance. VQ-Rec converts item text to a series of codes. Then with a lookup over a code embedding table, VQ-Rec can obtain item embeddings. Based on representation schema, VQ-Rec leverages contrastive learning among the next ground-truth items and negative items (i.e., semi-synthetic negative items and mixed-domain negative items). Finally, based on the code-embedding method, VQ-Rec implements a novel cross-domain fine-tuning method. TransRec [131] is pre-trained in one domain, and then is fine-tuned in another domain. The TransRec [131] achieves excellent recommendation performance. It is because that TransRec [131] first obtains knowledge from the source domain, and then transfers the obtained knowledge to the target domain. The three models mentioned above share a similar approach: they are initially pre-trained in one domain and then fine-tuned in another domain. IDA-SR [132] adopts a pre-trained text encoder to learn item representations by encoding item text descriptions. The pre-training tasks consist of next-item prediction, masked item prediction and permuted item prediction. These item representations can then be used for personalized recommendations through semantic transfer. SASRecV [133] leverages videos to denote an item. Compared with SASRec using an item ID to denote an item, SASRecV [133] can achieve comparable performance using an end-to-end framework. RECFORMER [134] mainly focuses on embedding initialization. By adopting behavior-tuned pre-trained language models for the initialization of ID-based SR models, RECFORMER [134] can achieve better recommendation performance without introducing additional computational costs. SRT-1K [135] calculates item embeddings using a trainable feature extractor, which can make the number of parameters be independent of the number of items. Additionally, SRT-1K [135] leverages contrastive learning to enhance catalog diversity representation. RECFORMER [136] flattens item attributes into sentences to represent items. The SR models mentioned above achieve improved recommendation performance by combining item IDs and multi-modal features.
@@
ODMT [130] fuses different modal information (i.e., text and images) together. Meanwhile, ODMT uses knowledge distillation to model weights for online recommendations. MMSR [121] integrates multi-modal information into nodes. MMMLP [137] incorporates multi-modal data into MLPs to enhance recommendation performance. p-RNN [138] leverages four different frameworks to combine some item features (i.e., images and text) and item IDs for modeling interaction sequences. Based on the aforementioned modal-based SR models, the following conclusions can be drawn:
@@
If we leverage a pre-trained modality model to extract features, we can obtain better performance than than training a model from scratch.
@@
Modal-based SR models can alleviate the cold-start problem.
@@
The hyper-parameters need to be searched carefully, as improper hyper-parameters easily lead to model collapse.
@@ 6.2 Generative Recommendation
Using semantic IDs prevents the embedding table size from increasing linearly with the number of items in datasets. It is because an item is denoted by a series of tokens instead of a unique item ID. The total number of tokens is far smaller than the number of all item IDs.
@@
In recent years, some generative sequential recommender models have been proposed. TIGER [139] is the first model that leverages semantic IDs for generative recommendation. MBGen [140] considers the role of behavioral information in generative recommendation. To consider the complementary nature between behavioral information and semantic information, EAGER [142] leverages a shared encoder and two separate decoders to decode the two kinds of information. CoST [143] leverages a contrastive quantization-based semantic tokenization approach to capture neighborhood relationships among items. LETTER [141] leverages semantic, collaborative and diversity regularization. The first two types of regularization can extract semantic information and collaborative signals, respectively. The diversity regularization can mitigate code assignment bias. To consider the multi-modal features of items, MMGRec [144] leverages a hierarchical quantization method (i.e., Graph RQ-VAE) by incorporating multi-modal features. Compared with traditional SR models, generative recommender models have some advantages as follows:
@@ 6.3.1 Large Language Models
PaLM: PaLM (Pathways Language Model) is developed by Google, which is one of the largest language models with up to about 540 billion parameters. It shows strong capabilities in multilingual tasks and source code generation.
@@ 6.3.2 LLMs for Recommendation
The third type of models leverages LLMs to generate high-quality data which is then combined with the original data to train SR models. The data generation process is formalized as follows:
@@
The most straightforward approach is to directly leverage LLMs for recommending personalized items. Given user ’s historical sequence and a target item, LLMs predict ’s preferences for the target item. Meanwhile, given the interacted item sequences, LLMs can generate corresponding recommendation lists. For example, P5 [145] can generate personalized recommend items by inputting historical interaction items into LLMs. There are different tasks, such as generating an item or a list of items, rating items, and ranking the order of the interacted items. GenRec [9] leverages an adaptation method (LoRA) to fine-tune and perform inference tasks on LLaMA. However, generative recommendation models have some weaknesses, such as the potential to generate non-existent items.
@@ 6.3.3 LLMs for Learning Semantic Embeddings
Some SR models use LLMs to initialize item embeddings. Firstly, RecInterpreter [149] leverages a linear layer to map item embeddings from an item ID embedding space to a text embedding space. By providing task-specific textual prompts and projected hidden representations, LLaMA is encouraged to generate textual descriptions of items in interaction sequences. The second task aims to discover residual items by task-specific textual prompts and hidden representations. To solve the problem of long-tail items, LLM-ESR [157] leverages LLMs to generate semantic embeddings for enhancing the semantic information of items. Common methods are to combine item ID embeddings and generated semantic embeddings for achieving better recommendation performance. To combine the advantages of traditional SR models and LLMs, E4SRec [152] uses item IDs as input. This ensures that all the generated recommendation items fall within candidate lists. Therefore, E4SRec can handle item IDs more effectively. Though knowledge can be transferred from some source domains to a target domain, the cold-start problem cannot be completely solved. LLMs cannot capture collaborative information effectively. Traditional cross-domain sequential recommendation models can capture collaborative signals while LLMs-based models excel at capturing semantic information. URLLM [153] incorporates both advantages. Firstly, URLLM leverages a dual-graph neural network to capture collaborative signals and transfer knowledge from a source domain to a target domain. Meanwhile, URLLM leverages a retrieve-generation model to incorporate structure information into an LLMs-based recommender model. Finally, URLLM combines collaborative information and semantic information to recommend items. From these models, we can draw the following conclusions: 1) Initializing item embeddings with a large language model can obtain recommendation performance gains. 2) Fine-tuning LLMs for recommendation tasks enables LLMs to learn domain-specific and domain-shared knowledge.
@@
Knowledge distillation and contrastive learning are effective techniques for extracting knowledge from LLMs. Knowledge distillation can enhance the efficiency of LLMs. SLIM [147] can obtain semantic knowledge from LLMs through knowledge distillation. DLLM2Rec [151] distills the knowledge from LLMs-based SR models to conventional SR models by leveraging two kinds of distillation (i.e., collaborative embedding distillation and importance-aware ranking distillation). LSVCR [156] is trained in two stages (i.e., personalized preferences alignment and recommendation-oriented fine-tuning). It uses contrastive learning between preferences generated by LLMs-based recommendation models and preferences generated by conventional SR models. Some conventional SR models can obtain richer semantic information.
LLMs often struggle with processing long text descriptions. LLM-TRSR [148] can leverage too-long text descriptions as the input of LLMs by segmenting original users’ interaction sequences. The LLMs-based SR models will consume lots of computational resources when processing long interacted sequences. To solve this problem, BAHE [162] leverages a behavior-aggregated hierarchical encoding method to enhance computational efficiency. By encoding a user’s separate interacted sequences in parallel, BAHE can obtain the user’s final representations by concatenating them together. Because LoRA’s shared parameters across all users limit its ability to recommend personalized items. Meanwhile, ultra-long interaction sequences might affect the efficiency of LLMs. RecLoRA [161] proposes a personalized LoRA module to overcome this limitation and designs a long-short modality retriever to retrieve interaction sequences of different lengths based on different modalities.

### v2

@@ 5.4.1 Common Features Fusion
In this section, we mainly introduce some SR models that fuse some common features (i.e., position, category, brand, etc.) and item IDs. The structure of these SR models is feature-agnostic to some extent. ICAI-SR [3] leverages a heterogeneous graph to capture complex item-to-attribute relationships by inner attribute aggregation and attribute-to-item aggregation, and then passes aggregation embeddings through an entity sequential model. MLP4Rec [117] designs a fusion mechanism to capture sequential, cross-channel and cross-feature correlations. SMLP4Rec [118] leverages layer normalization for a sequence mixing module, a feature mixing module and a channel mixing module simultaneously. Meanwhile, to enhance efficiency and effectiveness, SMLP4Rec leverages a parallel mode. RetaGNN [119] leverages a relational attentive GNN in which learnable weight matrices focus on various relations among users, items and attributes to have inductive and transferable properties. Then, RetaGNN [119] leverages Transformer to capture short-term and long-term temporal patterns. S3-Rec [120] leverages four self-supervised learning tasks (i.e., associated attribute prediction, masked item prediction, masked attribute prediction and segment prediction) to learn the associations between item IDs and item features.
In Transformer-based SR models, several advancements have been made in combining common features with item IDs. FDSA [121] first fuses different features by a vanilla attention mechanism. Then, FDSA leverages separate self-attention blocks to independently learn feature-level and item-level representations. Finally, FDSA fuses the two kinds of preferences for recommendation. SASRecF [25] is a sample model to fuse common feature embeddings and item ID embeddings through a concatenation operation. This fusion method is considered invasive [4], similar to an addition operation and a gating addition operation. The specific process is shown in part (a) of Figure 8. This fusion method might compromise the original information in item embeddings, potentially leading to negative effects. To address this issue, NOVA [4] leverages a non-invasive method to enhance the learning of attention matrices. In NOVA, key and query matrices are derived by integrating different feature embeddings and item embeddings while value matrices are directly obtained from item embeddings. However, NOVA [4] still leverages integrated embeddings to learn key and value matrices. The compounded embedding space might introduce random disturbances and share identical gradients, which might degrade recommendation performance. DIF-SR [25] leverages decoupled attention to adaptively learn different features by flexible gradients. MSSR [122] captures the associations among an item ID sequence and each feature sequence. ASIF [123] maintains semantic consistency between item IDs and features. Meanwhile, ASIF [123] makes full use of different kinds of features without compromising item embeddings, thanks to contrastive learning.
@@ 5.4.2 Context Information Fusion
Context information can reflect the specific environment where behaviors happen. SR models can provide more accurate recommendations by considering the context information. Bi-CASER [124] uses a context graph to represent context information, and then leverages bi-Transformer to learn users’ sequential behaviors. CRNNs [125] incorporates context information (e.g., interaction timestamps) by combining item embeddings with context embeddings to recommend personalized items. STAR [126] leverages stacked RNNs to model the sequential and context-aware information simultaneously. STEM [127] considers item behaviors and leverages a Transformer-based architecture for exploiting spatial and temporal information.
@@ 5.4.3 Special Features Fusion
In this section, we mainly introduce SR models designed to fuse special features. These SR models are typically applied to situations where certain special features (e.g., social networks and timestamps) are available. STEN [26] introduces temporal information and a social network to capture users’ fine-grained dynamic interests in an event-level direct paradigm. DUVRec [128] constructs a timespan-aware sequence graph and an attribute-augmented graph to learn item-view and factor-view user representations, respectively. Recommendation performance is then improved by fusing the two representations. Based on features and labels, STARec [129] retrieves a target user’s historical behaviors through a search-based retriever. Then, STARec [129] leverages a graph structure to enhance representation learning. SGCross [130] transfers knowledge from an auxiliary domain to a target domain from a personal view, a temporal view and a collaborative view. The knowledge encompasses personal preferences, dynamic preferences and collaborative preferences.
@@ 6.1 Multi-Modal SR
A user can be represented by the user’s profile. The profile may consist of user ’s demographic information or user ’s interacted items. An item can be denoted by an item ID embedding or a pre-trained modality embedding. Modality content features consist of text, images, videos, audio, text-image multi-modal pairs, and so on. ID-agnostic representations can be learned using pre-trained modality models. A common text encoder is BERT, which is usually used to extract semantic features from text. Swin Transformer is an image encoder, which can be leveraged to extract features from images. As shown in Figure 9. unlike ID-based SR models, one type of modal-based SR models leverages a single modality feature (e.g., text or images) to denote items. The other type of modal-based SR models leverages some modal features (e.g., text and images) to represent items [131]. In order words, the representations of item are obtained from the multi-modal features of item in modal-based SR models, which is formalized as follows:
@@
There are some single-modality-based SR models. UniSRec [8] leverages an MoE-enhanced adaptor and parametric whitening for domain adaptation. To learn universal sequence representations better, UniSRec proposes a seq-item contrastive task and a seq-to-seq contrastive task. The model is pre-trained in one domain and subsequently fine-tuned in another domain using either an inductive or a transductive approach. Meanwhile, UniSRec adopts two augmentation strategies (i.e., item drop and word drop). VQ-Rec [27] is a novel approach to learn vector-quantized item representations for solving the problem about over-emphasizing the effect of text features, which might lead to poor recommendation performance. VQ-Rec converts item text to a series of codes. Then with a lookup over a code embedding table, VQ-Rec can obtain item embeddings. Based on representation schema, VQ-Rec leverages contrastive learning among the next ground-truth items and negative items (i.e., semi-synthetic negative items and mixed-domain negative items). Finally, based on the code-embedding method, VQ-Rec implements a novel cross-domain fine-tuning method. TransRec [132] is pre-trained in one domain, and then is fine-tuned in another domain. TransRec [132] achieves excellent recommendation performance. It is because that TransRec [132] first obtains knowledge from a source domain, and then transfers the obtained knowledge to a target domain. The three models mentioned above share a similar approach: they are initially pre-trained in one domain and then fine-tuned in another domain. IDA-SR [133] adopts a pre-trained text encoder to learn item representations by encoding item text descriptions. The pre-training tasks consist of next-item prediction, masked item prediction and permuted item prediction. These item representations can then be used for personalized recommendations through semantic transfer. SASRecV [134] leverages videos to denote an item. Compared with SASRec using an item ID to denote an item, SASRecV [134] can achieve comparable performance using an end-to-end framework. RECFORMER [135] mainly focuses on embedding initialization. By adopting behavior-tuned pre-trained language models for the initialization of ID-based SR models, RECFORMER [135] can achieve better recommendation performance without introducing additional computational costs. SRT-1K [136] calculates item embeddings using a trainable feature extractor, which can make the number of parameters be independent of the number of items. Additionally, SRT-1K [136] leverages contrastive learning to enhance catalog diversity representation. RECFORMER [137] flattens item attributes into sentences to represent items. The SR models mentioned above achieve improved recommendation performance by combining item IDs and multi-modal features.
@@
There are some multi-modality-based SR models. ODMT [131] fuses different modal information (i.e., text and images) together. Meanwhile, ODMT uses knowledge distillation to model weights for online recommendations. MISSRec[138] leverages a multi-modal pre-training and transfer learning framework to integrate multi-modal information into SR models. MMSR [122] integrates multi-modal information into nodes. MMMLP [139] incorporates multi-modal data into MLPs to enhance recommendation performance. p-RNN [140] leverages four different frameworks to combine some item features (i.e., images and text) and item IDs for modeling interaction sequences. Based on the aforementioned modal-based SR models, the following conclusions can be drawn:
@@
If we leverage a pre-trained modality model to extract features, we can obtain better performance than training a model from scratch.
@@
Modal-based SR models can alleviate the cold-start problem to some extent.
@@
Hyper-parameters need to be searched carefully, as improper hyper-parameters easily lead to model collapse.
@@ 6.2 Generative Recommendation
Using semantic IDs prevents the embedding table size from increasing linearly with the number of items in a data. It is because an item is denoted by a series of tokens instead of a unique item ID. The total number of tokens is far smaller than the number of all item IDs.
@@
In recent years, some generative sequential recommender models have been proposed. TIGER [141] is the first model that leverages semantic IDs for generative recommendation. MBGen [142] considers the role of behavioral information in generative recommendation. To consider the complementary nature between behavioral information and semantic information, EAGER [144] leverages a shared encoder and two separate decoders to decode the two kinds of information. CoST [145] leverages a contrastive quantization-based semantic tokenization approach to capture neighborhood relationships among items. LETTER [143] leverages semantic, collaborative and diversity regularization. The first two types of regularization can extract semantic information and collaborative signals, respectively. The diversity regularization can mitigate code assignment bias. To consider the multi-modal features of items, MMGRec [146] leverages a hierarchical quantization method (i.e., Graph RQ-VAE) by incorporating multi-modal features. SpecGR [147] leverages a drafter model to generate some candidate items, including some existing items and new items. Meanwhile, SpecGR [147] leverages a verifier to decide whether to accept or reject candidate items, enabling SpecGR to recommend new items. Compared with traditional SR models, generative recommender models have some advantages as follows:
@@ 6.3.1 Large Language Models
PaLM: PaLM (Pathways Language Model) is developed by Google, which is one of the largest language models. It shows strong capabilities in multilingual tasks and source code generation.
@@
Baichuan: Baichuan contains a series of large language models developed by Baichuan Intelligent Technology, including Baichuan 2 and Baichuan-M1.
DeepSeek: DeepSeek 777https://github.com/deepseek-ai has recently achieved remarkably impressive performance on various metrics in comparison with the most powerful LLMs.
@@ 6.3.2 LLMs for Recommendation
As shown in part (c) of Figure 12, the third type of models leverages LLMs to generate high-quality data which is then combined with the original data to train SR models. The data generation process is formalized as follows:
@@
The most straightforward approach is to directly leverage LLMs for recommending personalized items. Given user ’s historical sequence and a target item, LLMs predict ’s preferences for the target item. Meanwhile, given the interacted item sequences, LLMs can generate corresponding recommendation lists. For example, P5 [148] can generate personalized recommend items by inputting historical interaction items into LLMs. There are different tasks, such as generating an item or a list of items, rating items, and ranking the order of the interacted items. POD[150] converts discrete prompt into a set of continuous prompt vectors. This method can reduce the inference time by bridging the gap between IDs and words. GenRec [9] leverages an adaptation method (LoRA) to fine-tune and perform inference tasks on LLaMA. However, these models have some weaknesses, such as the potential to generate non-existent items.
@@ 6.3.3 LLMs for Learning Semantic Embeddings
Some SR models use LLMs to initialize item embeddings. Firstly, RecInterpreter [153] leverages a linear layer to map item embeddings from an item ID embedding space to a text embedding space. By providing task-specific textual prompts and projected hidden representations, LLaMA is encouraged to generate textual descriptions of items in interaction sequences. The second task aims to discover residual items by task-specific textual prompts and hidden representations. To solve the problem of long-tail items, LLM-ESR [164] leverages LLMs to generate semantic embeddings for enhancing the semantic information of items. Common methods are to combine item ID embeddings and generated semantic embeddings for achieving better recommendation performance. To combine the advantages of traditional SR models and LLMs, E4SRec [156] uses item IDs as input. This method ensures that all the generated and recommended items fall within candidate lists. Therefore, E4SRec can handle item IDs more effectively. LLMEmb [162] is a novel method that leverages LLMs to generate item embeddings. To integrate collaborative signals into LLM-generated embeddings, LLMEmb leverages a recommendation adaptation training strategy. Though knowledge can be transferred from some source domains to a target domain, the cold-start problem cannot be completely solved. LLMs cannot capture collaborative information effectively. Traditional cross-domain sequential recommendation models can capture collaborative signals while LLMs-based models excel at capturing semantic information. URLLM [157] incorporates both advantages. Firstly, URLLM leverages a dual-graph neural network to capture collaborative signals and transfer knowledge from a source domain to a target domain. Meanwhile, URLLM leverages a retrieve-generation model to incorporate structure information into an LLMs-based recommender model. Finally, URLLM combines collaborative information and semantic information to recommend items. LLMCDSR [170] leverages LLMs to predict unobserved cross-domain interactions, which improves the recommendation performance for non-overlapped users. LLMCDSR [170] leverages a collaborative-textual contrastive pre-training approach to integrate collaborative information into textual features. From these models, we can draw the following conclusions: 1) Initializing item embeddings with a large language model can obtain recommendation performance gains. 2) Fine-tuning LLMs for recommendation tasks enables LLMs to learn domain-specific and domain-shared knowledge.
@@
Knowledge distillation and contrastive learning are effective techniques for extracting knowledge from LLMs. Knowledge distillation can enhance the efficiency of LLMs. SLIM [151] can obtain semantic knowledge from LLMs through knowledge distillation. DLLM2Rec [155] distills the knowledge from LLMs-based SR models to conventional SR models by leveraging two kinds of distillation (i.e., collaborative embedding distillation and importance-aware ranking distillation). LSVCR [163] is trained in two stages (i.e., personalized preferences alignment and recommendation-oriented fine-tuning). It uses contrastive learning between preferences generated by LLMs-based recommendation models and preferences generated by conventional SR models. Therefore, these conventional SR models can obtain richer semantic information. SLMRec[161] leverages a layer-wise knowledge distillation approach to enhance the efficiency of SR models based on LLMs.
LLMs often struggle with processing long text descriptions. LLM-TRSR [152] can leverage too-long text descriptions as the input of LLMs by segmenting original users’ interaction sequences. The LLMs-based SR models will consume lots of computational resources when processing long interacted sequences. To solve this problem, BAHE [169] leverages a behavior-aggregated hierarchical encoding method to enhance computational efficiency. By encoding a user’s separate interacted sequences in parallel, BAHE can obtain the user’s final representations by concatenating them together. Because LoRA’s shared parameters across all users limit its ability to recommend personalized items. Meanwhile, ultra-long interaction sequences might affect the efficiency of LLMs. Therefore, RecLoRA [168] proposes a personalized LoRA module to overcome this limitation and designs a long-short modality retriever to retrieve interaction sequences of different lengths based on different modalities.